D) Find the interest rate or performance rate for each of the following: 1) You borrow $900 and promise to pay $972 at the end of a year;

2) You borrow $75.000 and promise to pay $280.500 at the end of 15 years.

Please explain each step
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1.

972=900*(1+r/100)

(972/900)=(1+r/100)

(1+r/100)=1.08

r=(1.08-1)*100

=8%

2.

280.5=75*(1+r/100)

(280.5/75)^(1/15)=(1+r/100)

(1+r/100)=1.0919

r=1.0919-1

=9.19%(Approx).